OpenStudy (ammarah):

Tell whether the given side lengths of a triangle can represent a right triangle.

OpenStudy (ammarah):

9,10,and 15

OpenStudy (ammarah):

6,10,and 2 to the sqrt 34

OpenStudy (ammarah):

@mathstudent55

OpenStudy (kuoministers):

i would say no to the first set of numbers

OpenStudy (ammarah):

yes cuz it would be obtuse?

OpenStudy (kuoministers):

You can try putting the numbers into the Pythagorean theorem and see if it works. Or, you can reduce the numbers by their greatest common factor and see if it reduces down to one of the common right triangles

OpenStudy (kuoministers):

probably would be :)

OpenStudy (ammarah):

but i dont get the next question with the sqrt

OpenStudy (kuoministers):

6,10,and 2 to the sqrt 34 this set would be a right angle triangle

OpenStudy (kuoministers):

\[\sqrt{x}^{2}\] would be??

OpenStudy (ammarah):

?

OpenStudy (kuoministers):

\[(2\sqrt{34} )^{2}\]

OpenStudy (ammarah):

yes so is that the longest side?

OpenStudy (ammarah):

how do u know when it is the longest side?

OpenStudy (kuoministers):

it would be x because a sqrt cancels out a square

OpenStudy (kuoministers):

its like saying what is the longest? 2 4 6.. you just know!!! square them all and see which is the longest side

OpenStudy (ammarah):

6...

OpenStudy (ammarah):

wait so it is the longest side..

OpenStudy (mathstudent55):

Try the numbers in the Pythagorean Theorem. Always use the largest number as the hypotenuse. If it makes it true, then it is a right triangle. First example: 9, 10, 15 Is 9^2 + 10^2 = 15^2? Is 81 + 100 = 225? Is 181 = 225? Answer is no, so 9, 10, 15 do not represnt the sides of a right triangle.

OpenStudy (kuoministers):

yea it is the longest side

OpenStudy (mathstudent55):



OpenStudy (ammarah):

yes

OpenStudy (mathstudent55):

If so, that is 6, 10 and 2 times square root of 34, not 2 to the sqrt 34.

OpenStudy (ammarah):

yea yeah thats wahat i meant..

OpenStudy (mathstudent55):

Try again just like before. First, if you have a calculator, make sure 2*sqrt(34) is larger than 10. I checked and it is. Is 6^2 + 10^2 = (2*sqrt(34))^2 ?

OpenStudy (mathstudent55):

6^2 + 10^2 = (2*sqrt(34))^2 36 + 100 = 4*34 136 = 136 Since this is true, that means 6, 10, and 2*sqtr(34) are the sides of a right triangle.

OpenStudy (ammarah):

(2*sqrt(34) this is like 5. something..

OpenStudy (kuoministers):

not really

OpenStudy (ammarah):

huh....than what is it? cuz im typing it in my calculator...

OpenStudy (kuoministers):

i would be 11.66......?

OpenStudy (kuoministers):

it*

OpenStudy (ammarah):

how did u get that?

OpenStudy (ammarah):

im typing it but.....it sais 5.something

OpenStudy (mathstudent55):

Square root of 34 is 5.83..., 2 * sqrt(34) = 11.66... which is greater than 10

OpenStudy (ammarah):

OH!

OpenStudy (ammarah):

now i get it..

OpenStudy (ammarah):

I HAVE ALOT OF OTHER QUESTIONS PLEASE HELP ILL TAG U GUYS!! K??

OpenStudy (kuoministers):

sure if i get a medal :D

OpenStudy (kuoministers):

haha thanks

OpenStudy (mathstudent55):

yes, tag me